OpenStudy (madeleine2):

Which phrase describes the end behavior of the graph of ƒ(x) = –x^5 – 4x^4 + 7x^2 – 2? a. up to the right, down to the left b. down to the right, up to the left c. up in both directions d. down in both directions

OpenStudy (mathmale):

Let me point out that this is a 5th degree (odd degree) polynomial, and that the sign of the dominant term is negative. y=x^3 has the simplest curve other than y=x (which is just a straight line). In which quadrant does y=x^3 begin? End? Now supposing you had y=-x^3. How would your graph be different? In which quadrant would it begin and in which would it end, as you move from left to right? Your answers will also apply to y=-x^5.

OpenStudy (mathmale):

Glad to help, but need your involvement. if you've already moved beyond this problem, please close this post.